# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v2

@@ 4.1 The case of general parameters
$P_{0}(N)=(N+4)^{2}(N+6)^{2},\>\>P_{1}(N)=-375+400N+843N^{2}+322N^{3}+35N^{4}$
$P_{2}(N)=(45+622N+259N^{2})(N-1)^{2},\>\>P_{3}(N)=-255(N-2)^{2}(N-1)^{2}.$ (4.15)
From the explicit forms (4.14), (4.8) and (4.2) one sees an obvious pattern in that the degree of the polynomial coefficients within each case is the same, and equal to l1l-1. However the cancellations which take place in going from the difference equation in Proposition 4.4 to the one for $\{T_{l}(N)\}$ makes this difficult to anticipate.
@@
1. We know that the matrix differential equation of Proposition 3.2 has a unique solution with $d_{p,0}=\delta_{p,l}\,(p=0,\dots,l)$ in (3.11); recall Remark 3.1 point 1. As a consequence it must be that the scalar differential equation (4.1) has a unique power series solution subject only the the normalisation $I_{n}(0)=1$ (i.e. that the constant term in the power series is fixed). For this latter property to hold, it must be that the corresponding difference equation (4.10) (corrected according to Proposition 4.3) must have the property that the coefficients of $T_{l}(-1),T_{l}(-2),\dots,T_{l}(-[(l-1)/2])$, whenever they appear, must vanish. Note that this property is illustrated in (4.8) and (4.2).
2. In relation to computing the power series solution for general $l$, there is no advantage in first reducing the matrix differential equation down to a scalar differential equation, rather use of the analogue of (3.18) (which is based on the matrix differential equation) is more efficient. As already mentioned in Remark 3.1, point 2, a quantification of the complexity when using this approach is given in Appendix B.
@@
$\bigg(x^{2}{d^{3}\over dx^{3}}+10x{d^{2}\over dx^{2}}+(20-4x){d\over dx}-10\bigg)y(x)=0,$
$\bigg(x^{3}{d^{4}\over dx^{4}}+28x^{2}{d^{3}\over dx^{3}}+(230-10x)x{d^{2}\over dx^{2}}+(540-122x){d\over dx}+(-270+9x)\bigg)y(x)=0$
$\bigg(x^{4}{d^{5}\over dx^{5}}+60x^{3}{d^{4}\over dx^{4}}+(1238-20x)x^{2}{d^{3}\over dx^{3}}+(10268-662x)x{d^{2}\over dx^{2}}+(28288-6108x+64x^{2}){d\over dx}$
${}\qquad\qquad+(-14144+640x)\bigg)y(x)=0$
$\bigg(x^{5}{d^{6}\over dx^{6}}+110x^{4}{d^{5}\over dx^{5}}+(4538-35x)x^{3}{d^{4}\over dx^{4}}+(87008-2422x)x^{2}{d^{3}\over dx^{3}}+(765908-55800x+$
${}\quad+259x^{2})x{d^{2}\over dx^{2}}+(2431000-485408x+8392x^{2}){d\over dx}+(-1215500+59300x-225x^{2})\bigg)y(x)=0,$
@@
As revised in Appendix A, the work [39] has given an alternative way to derive the degree $l+1$ scalar linear differential equation satisfied by the matrix integral (1.7). In the case $l=2$ this is made explicit in [39, Eq. (97)], and consistency with the third order equation listed above can be checked. Implementing the recursive strategy leading to (A.12) of the Appendix in the cases $l=3,4,5$ using computer algebra establishes consistency of the results obtained by the alternative method for all the differential equations listed in Proposition 4.5.
@@
$v(s):=-s{d\over ds}\log\bigg(s^{l^{2}/2}e^{-s/4}\Big\langle(\det U)^{l}e^{{1\over 2}\sqrt{s}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}\bigg);$ (4.16)
@@
$(sv^{\prime\prime})^{2}+v^{\prime}(4v^{\prime}-1)(v-sv^{\prime})-{l^{2}\over 16}=0,$ (4.17)
@@
$v(s)=-l^{2}+{s\over 8}+{\rm O}(s^{2}).$ (4.18)
It was shown in [33] that this allowed for the generation of the first $l$ even powers recursively, as was the need of the application therein. However, missed in [33] was the fact made explicit in [39] that the boundary condition (4.18) does not uniquely determine a solution of (4.18), with there in fact being a family of solutions which require the specification of the coefficient of $s^{2l+1}$ to be made unique beyond degree $s^{2l}$, and which was incorrectly set to zero (albeit without consequence to the application given their related to the moments of the first deviative of the Riemann zeta function).
@@
Acknowledgement
[…]
Appendix A: Alternative derivations from [39]
@@ Acknowledgement
Definition 1.
Let $\alpha\in\mathbb{R}_{\geq 0}$. Let $l\geq 1,h\geq 0$ be integers. Let $Y=(m_{1},\ldots,m_{h})$ be a Young diagram. Define
$\tau_{l,Y,\alpha}(x):=\det\!\big[I_{j+k+\alpha+m_{l-k}}(2\sqrt{x})\big]_{j,k=0}^{\,l-1}.$
Here we adopt the convention that $m_{h+1}=\cdots=m_{l}=0$ whenever $h<l$. When $Y=\emptyset$, we simply write $\tau_{l,\emptyset,\alpha}$ as $\tau_{l,\alpha}$ for short.
In particular, when $Y=\emptyset$ and $\alpha=1$, the above definition reduces to $\tau_{l}(x)$ as given in (A.1). Using the trace of a square matrix, we define the following operations.
Definition 2.
@@
$(T_{n}\tau_{l,Y,\alpha})(x):=\mathrm{Tr}\left(\mathrm{adj}\big(\big[I_{j+k+\alpha+m_{l-k}}(2\sqrt{x}\big)\big]_{j,k=0}^{\,l-1})\big[I_{j+k+\alpha+m_{l-k}+n}(2\sqrt{x})\big]_{j,k=0}^{\,l-1}\right)$
@@
$(S_{n}\tau_{l,Y,\alpha})(x):$
$=\mathrm{Tr}\left(\mathrm{adj}\big(\big[I_{j+k+\alpha+m_{l-k}}(2\sqrt{x})\big]_{j,k=0}^{\,l-1}\big)\big[(j+k+\alpha+m_{l-k}+n)I_{j+k+\alpha+m_{l-k}+n}(2\sqrt{x})\big]_{j,k=0}^{\,l-1}\right),$
@@
In deriving the higher order linear differential equation for $\tau_{l,\alpha}$, we mainly use a special form of Young diagram, namely
@@
of weight $m$, where $m\geq j\geq 1$. By applying [39, Thm. 3.1] to general Hankel determinants shifted by $Y_{m,j}$ and specializing to $\tau_{l,Y_{m,j},\alpha}$, we obtain the following inverse linear system.
@@
where $\mathbf{B}_{m}=(b_{j,k})_{i,j=1,\ldots,m}$ is the $m\times m$-matrix satisfying
$b_{j,k}=\begin{cases}(-1)^{j+k-1}/k(k+1)&{\color[rgb]{0,0,0}j\leq k\leq m-1};\\ -1/j&k=j-1;\\ 0&k<j-1;\\ (-1)^{j-1}/m&k=m.\end{cases}.$ (A.3)
@@
$\frac{d}{dx}I_{\beta}(2\sqrt{x})=\frac{I_{\beta+1}(2\sqrt{x})}{\sqrt{x}}+\frac{\beta}{2x}I_{\beta}(2\sqrt{x}),\quad I_{\beta+2}(2\sqrt{x})=I_{\beta}(2\sqrt{x})-\frac{\beta+1}{\sqrt{x}}I_{\beta+1}(2\sqrt{x}),$ (A.4)
we can express $\tau_{l,Y_{1,1},\alpha}$, $\tau_{l,Y_{2,1},\alpha}$, and $\tau_{l,Y_{2,2},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. The following lemmas for general $\alpha$ are obtained by a similar argument to that used in [39, Prop. 4.1 and Prop. 4.2] for $\alpha=1$.
Lemma 1.
@@
$T_{1}\tau_{l,Y,\alpha}(x)=\sqrt{x}\frac{d}{dx}\tau_{l,Y,\alpha}(x)-\frac{l^{2}+m+l(\alpha-1)}{2\sqrt{x}}\tau_{l,Y,\alpha}.$
In particular, if $Y=\emptyset$, then
$\tau_{l,Y_{1,1},\alpha}(x)=\sqrt{x}\frac{d}{dx}\tau_{l,\alpha}(x)-\frac{l^{2}+l(\alpha-1)}{2\sqrt{x}}\tau_{l,\alpha}.$ (A.5)
Lemma 2.
@@
$\tau_{l,Y_{2,1},\alpha}=$ $\frac{x}{2}\frac{d^{2}}{dx^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha+1)+\alpha-1}{2}\frac{d}{dx}\tau_{l,\alpha}$
$+\frac{(l^{2}+l(\alpha-1))(l^{2}+l(\alpha+3)+2\alpha)+4lx}{8x}\tau_{l,\alpha},$ (A.6)
$\tau_{l,Y_{2,2},\alpha}=$ $\frac{x}{2}\frac{d^{2}}{dx^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha-3)-(\alpha-1)}{2}\frac{d}{dx}\tau_{l,\alpha}$
$+\frac{(l^{2}+l(\alpha-1))(l^{2}+l(\alpha-5)+4-2\alpha)-4lx}{8x}\tau_{l,\alpha}.$ (A.7)
For a general weight $m$ of Young diagrams, we can recursively obtain an expression for $\tau_{l,Y_{m,j},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. Thus, by identity (A.4), we can rewrite operations involving $T_{h}$ in Proposition A.1 in terms of $S_{h-1}$ and $T_{h-2}$. Combining Lemmas 1 and 2, we then obtain the following formula for $\tau_{l,Y_{m,j},\alpha}$, which depends recursively on $m$.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\end{pmatrix}$= $-\,\sqrt{x}\mathbf{B}_{m}\begin{pmatrix}\frac{d}{dx}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \frac{d}{dx}\tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}+\frac{l^{2}+l(\alpha-1)+m-1}{2\sqrt{x}}\mathbf{B}_{m}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}$
@@
where $\mathbf{B}_{m}$ is given as in (A.3), $\mathbf{C}^{(1)}_{m}=(c_{j,k}^{(1)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-1\end{subarray}}$ is an $m\times(m-1)$-matrix satisfying
$c_{j,k}^{(1)}=\begin{cases}(-1)^{j+k}\big(\frac{m-1-k}{k+1}+\frac{2l-k+\alpha}{k(k+1)}\big)&\text{if }j\leq k\leq m-1;\\ -\frac{1}{k+1}\big(k(2l-k+\alpha)+m-1-k\big)&\text{if }k=j-1;\\ 0&\text{if }k<j-1,\end{cases},$ (A.9)
and $\mathbf{C}^{(2)}_{m}=(c_{j,k}^{(2)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-2\end{subarray}}$ is an $m\times(m-2)$-matrix satisfying
$c_{j,k}^{(2)}=\begin{cases}(-1)^{j+k}\frac{l+2}{(k+1)(k+2)}&j-1\leq k\leq m-2;\\ \frac{j-l-2}{j}&k=j-2;\\ 0&k<j-2.\end{cases}.$ (A.10)
@@
The above proposition generalizes [39, Prop. 4.3] from the case $\alpha=1$ to general $\alpha$. It can be rewritten as a first order matrix linear differential equation for the vector on the LHS analogous to (but distinct from) that implied by Proposition 3.2, with $n=m$, $\beta=2$ and $B_{p}$ given by (3.16) with $q=m+\alpha$. On this latter point, note that the entries in the corresponding vectors agree in the final entry but involve different functions otherwise.
Theorem 1.
@@
$\tau_{l,Y_{l+1,l+1},\alpha}(x)=x^{-\frac{l+1}{2}}\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(\alpha)\,x^{j}\right)\frac{d^{n}\tau_{l,\alpha}(x)}{dx^{n}},$ (A.11)
where the coefficients $b_{j,n}^{(l+1)}(\alpha)$ are given recursively. Moreover, we have the following equation
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(\alpha)\,x^{j}\right)\frac{d^{n}\tau_{l,\alpha}(x)}{dx^{n}}=0.$ (A.12)
@@
$\tau_{l,Y_{l+1,l+1},\alpha}(x)=$ $\frac{1}{l+1}\left(\sqrt{x}\frac{d\tau_{l,Y_{l,l},\alpha}(x)}{dx}-\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)\right)+\frac{l(l+\alpha)}{(l+1)\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)$
@@
$\tau_{l,Y_{l-1,l-1},\alpha}(x)=\mathrm{Tr}\left(\mathrm{adj}\big(\big[I_{j+k+\alpha+1}(2\sqrt{x}\big)\big]_{j,k=0}^{\,l-1})\big[I_{j+k+\alpha}(2\sqrt{x})\big]_{j,k=0}^{\,l-1}\right).$
By the observation $\tau_{l,Y_{l,l},\alpha}(x)=\det(I_{j+k+\alpha+1}(2\sqrt{x}))_{j,k=0,\ldots,l-1}$ and the recurrence relation
$I_{j+k+\alpha}(2\sqrt{x})=\sqrt{x}\frac{d}{dx}I_{j+k+\alpha+1}(2\sqrt{x})+\frac{\alpha+j+k+1}{2\sqrt{x}}I_{j+k+\alpha+1}(2\sqrt{x}),$
@@
$\sqrt{x}\frac{d}{dx}\tau_{l,Y_{l,l},\alpha}(x)+\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)-\tau_{l,Y_{l-1,l-1},\alpha}(x)=0.$
Then by (A.13), we have the identity
$\tau_{l,Y_{l+1,l+1},\alpha}\equiv 0.$
Hence, using above and (A.11), we obtain (A.12). ∎
@@
As a consequence of Theorem 1 with $\alpha=1$, we obtain the following result, which provides a method to derive the degree $l+1$ linear differential equation satisfied by (1.7).
@@
Let $l\geq 1$. Let $b_{j,n}^{(l+1)}(1)$ be given as in Theorem 1 with $\alpha=1$. Then
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(1)\,x^{j}\right)\frac{d^{n}\tau_{l}(x)}{dx^{n}}=0.$
Finally, applying Corollary A.1 and verifying with the code [^5] https://drive.google.com/file/d/1gd4kQO9ABKhALrzeYos-DbeLcelQmik6/view?usp=share_link, we re-obtain Proposition 4.5.
Appendix B: Computational aspects (authored by F. Bornemann)
B.4 $D$-finiteness

### v3

@@ 4.1 The case of general parameters
$P_{0}(N)=(N+4)^{2}(N+6)^{2},\qquad P_{1}(N)=-375+400N+843N^{2}+322N^{3}+35N^{4},$
$P_{2}(N)=\bigl(45+622N+259N^{2}\bigr)(N-1)^{2},\qquad P_{3}(N)=-255(N-2)^{2}(N-1)^{2}.$ (4.11)
From the explicit forms (4.10), (4.7) and (4.11), one sees an obvious pattern in that the degree of the polynomial coefficients within each case is the same, and equal to l1l-1. However, the cancellations which take place in going from the difference equation in Proposition 4.4 to the one for $\{T_{l}(N)\}$ makes this difficult to anticipate.
@@
  • (1)
    We know that the matrix differential equation of Proposition 3.3 has a unique solution with $d_{p,0}=\delta_{p,l}(p=0,\dots,l)$ in (3.8); recall Remark 3.4 point 1. As a consequence, it must be that the scalar differential equation (4.1) has a unique power series solution subject only the normalisation $I_{n}(0)=1$ (i.e., that the constant term in the power series is fixed). For this latter property to hold, it must be that the corresponding difference equation (4.9) (corrected according to Proposition 4.3) must have the property that the coefficients of $T_{l}(-1),T_{l}(-2),\dots,T_{l}(-[(l-1)/2])$}, whenever they appear, must vanish. Note that this property is illustrated in (4.7) and (4.11).
  • (2)
    In relation to computing the power series solution for general $l$, there is no advantage in first reducing the matrix differential equation down to a scalar differential equation, rather use of the analogue of (3.15) (which is based on the matrix differential equation) is more efficient. As already mentioned in Remark 3.4, point 2, a quantification of the complexity when using this approach is given in Appendix B.
@@
$\left(x^{2}{{\rm d}^{3}\over{\rm d}x^{3}}+10x{{\rm d}^{2}\over{\rm d}x^{2}}+(20-4x){{\rm d}\over{\rm d}x}-10\right)y(x)=0,$
$\left(x^{3}{{\rm d}^{4}\over{\rm d}x^{4}}+28x^{2}{{\rm d}^{3}\over{\rm d}x^{3}}+(230-10x)x{{\rm d}^{2}\over{\rm d}x^{2}}+(540-122x){{\rm d}\over{\rm d}x}+(-270+9x)\right)y(x)=0,$
$\left(x^{4}{{\rm d}^{5}\over{\rm d}x^{5}}+60x^{3}{{\rm d}^{4}\over{\rm d}x^{4}}+(1238-20x)x^{2}{{\rm d}^{3}\over{\rm d}x^{3}}+(10268-662x)x{{\rm d}^{2}\over{\rm d}x^{2}}\right$.
$\left.\qquad+\bigl(28288-6108x+64x^{2}\bigr){{\rm d}\over{\rm d}x}+(-14144+640x)\right)y(x)=0,$
$\left(x^{5}{{\rm d}^{6}\over{\rm d}x^{6}}+110x^{4}{{\rm d}^{5}\over{\rm d}x^{5}}+(4538-35x)x^{3}{{\rm d}^{4}\over{\rm d}x^{4}}+(87008-2422x)x^{2}{{\rm d}^{3}\over{\rm d}x^{3}}\right$.
$\qquad+\bigl(765908-55800x+259x^{2}\bigr)x{{\rm d}^{2}\over{\rm d}x^{2}}+\bigl(2431000-485408x+8392x^{2}\bigr){{\rm d}\over{\rm d}x}$
$\qquad+\left.\bigl(-1215500+59300x-225x^{2}\bigr)\right)y(x)=0,$
@@
As revised in Appendix A, the work [39] has given an alternative way to derive the degree l+1{l+1} scalar linear differential equation satisfied by the matrix integral (1.7). In the case $l=2$, this is made explicit in [39, equation (97)], and consistency with the third-order equation listed above can be checked. Implementing the recursive strategy leading to (A.10) of Appendix A, in the cases $l=3,4,5$ using computer algebra establishes consistency of the results obtained by the alternative method for all the differential equations listed in Proposition 4.6.
@@
$v(s):=-s{{\rm d}\over{\rm d}s}\log\big(s^{l^{2}/2}{\rm e}^{-s/4}\big\langle(\det U)^{l}{\rm e}^{{1\over 2}\sqrt{s}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}\big);$
@@
$(sv^{\prime\prime})^{2}+v^{\prime}(4v^{\prime}-1)(v-sv^{\prime})-{l^{2}\over 16}=0,$
@@
$v(s)=-l^{2}+{s\over 8}+{\rm O}\bigl(s^{2}\bigr).$ (4.12)
It was shown in [33] that this allowed for the generation of the first $l$ even powers recursively, as was the need of the application therein. However, missed in [33] was the fact made explicit in [39] that the boundary condition (4.12) does not uniquely determine a solution of (4.12), with there in fact being a family of solutions which require the specification of the coefficient of $s^{2l+1}$ to be made unique beyond degree $s^{2l}$, and which was incorrectly set to zero (albeit without consequence to the application given there related to the moments of the first derivative of the Riemann zeta function).
@@
Appendix A Alternative derivations from [39]
@@
Definition A.1.
Let $\alpha\in\mathbb{R}_{\geq 0}$. Let $l\geq 1,h\geq 0$ be integers. Let $Y=(m_{1},\ldots,m_{h})$ be a Young diagram. Define $\tau_{l,Y,\alpha}(x):=\det\!\big[I_{j+k+\alpha+m_{l-k}}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}$. Here we adopt the convention that $m_{h+1}=\cdots=m_{l}=0$ whenever $h<l$. When $Y=\varnothing$, we simply write $\tau_{l,\varnothing,\alpha}$ as $\tau_{l,\alpha}$ for short.
In particular, when $Y=\varnothing$ and $\alpha=1$, the above definition reduces to $\tau_{l}(x)$ as given in (A.1). Using the trace of a square matrix, we define the following operations.
Definition A.2.
@@
$(T_{n}\tau_{l,Y,\alpha})(x):=\operatorname{Tr}\left(\mathrm{adj}\bigl(\big[I_{j+k+\alpha+m_{l-k}}(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1})\big[I_{j+k+\alpha+m_{l-k}+n}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}\right)$
@@
$(S_{n}\tau_{l,Y,\alpha})(x):={}$ $\operatorname{Tr}\bigl(\mathrm{adj}\bigl(\big[I_{j+k+\alpha+m_{l-k}}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}\bigr)$
$\times\big[(j+k+\alpha+m_{l-k}+n)I_{j+k+\alpha+m_{l-k}+n}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}\bigr),$
@@
In deriving the higher-order linear differential equation for $\tau_{l,\alpha}$, we mainly use a special form of Young diagram, namely
@@
of weight $m$, where $m\geq j\geq 1$. By applying [39, Theorem 3.1] to general Hankel determinants shifted by $Y_{m,j}$ and specializing to $\tau_{l,Y_{m,j},\alpha}$, we obtain the following inverse linear system.
@@
where $\mathbf{B}_{m}=(b_{j,k})_{j,k=1,\ldots,m}$ is the $m\times m$ matrix satisfying
$b_{j,k}=\begin{cases}(-1)^{j+k-1}/k(k+1),&{\color[rgb]{0,0,0}j\leq k\leq m-1},\\ -1/j,&k=j-1,\\ 0,&k<j-1,\\ (-1)^{j-1}/m,&k=m.\end{cases}.$ (A.3)
@@
$\frac{\rm d}{{\rm d}x}I_{\beta}\bigl(2\sqrt{x}\bigr)=\frac{I_{\beta+1}\bigl(2\sqrt{x}\bigr)}{\sqrt{x}}+\frac{\beta}{2x}I_{\beta}\bigl(2\sqrt{x}\bigr),$
$I_{\beta+2}\bigl(2\sqrt{x}\bigr)=I_{\beta}\bigl(2\sqrt{x}\bigr)-\frac{\beta+1}{\sqrt{x}}I_{\beta+1}\bigl(2\sqrt{x}\bigr),$ (A.4)
we can express $\tau_{l,Y_{1,1},\alpha}$, $\tau_{l,Y_{2,1},\alpha}$, and $\tau_{l,Y_{2,2},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. The following lemmas for general $\alpha$ are obtained by a similar argument to that used in [39, Propositions 4.1 and 4.2] for $\alpha=1$.
Lemma A.4.
@@
$T_{1}\tau_{l,Y,\alpha}(x)=\sqrt{x}\frac{\rm d}{{\rm d}x}\tau_{l,Y,\alpha}(x)-\frac{l^{2}+m+l(\alpha-1)}{2\sqrt{x}}\tau_{l,Y,\alpha}.$
In particular, if $Y=\varnothing$, then
$\tau_{l,Y_{1,1},\alpha}(x)=\sqrt{x}\frac{\rm d}{{\rm d}x}\tau_{l,\alpha}(x)-\frac{l^{2}+l(\alpha-1)}{2\sqrt{x}}\tau_{l,\alpha}.$ (A.5)
Lemma A.5.
@@
$\tau_{l,Y_{2,1},\alpha}={}$ $\frac{x}{2}\frac{{\rm d}^{2}}{{\rm d}x^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha+1)+\alpha-1}{2}\frac{\rm d}{{\rm d}x}\tau_{l,\alpha}$
$+\frac{\bigl(l^{2}+l(\alpha-1)\bigr)\bigl(l^{2}+l(\alpha+3)+2\alpha\bigr)+4lx}{8x}\tau_{l,\alpha},$ (A.6)
$\tau_{l,Y_{2,2},\alpha}={}$ $\frac{x}{2}\frac{{\rm d}^{2}}{{\rm d}x^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha-3)-(\alpha-1)}{2}\frac{\rm d}{{\rm d}x}\tau_{l,\alpha}$
$+\frac{\bigl(l^{2}+l(\alpha-1)\bigr)\bigl(l^{2}+l(\alpha-5)+4-2\alpha\bigr)-4lx}{8x}\tau_{l,\alpha}.$ (A.7)
For a general weight $m$ of Young diagrams, we can recursively obtain an expression for $\tau_{l,Y_{m,j},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. Thus, by identity (A.4), we can rewrite operations involving $T_{h}$ in Proposition A.3 in terms of $S_{h-1}$ and $T_{h-2}$. Combining Lemmas A.4 and A.5, we then obtain the following formula for $\tau_{l,Y_{m,j},\alpha}$, which depends recursively on $m$.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\end{pmatrix}$={} $-\sqrt{x}\mathbf{B}_{m}\begin{pmatrix}\frac{\rm d}{{\rm d}x}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \frac{\rm d}{{\rm d}x}\tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}+\frac{l^{2}+l(\alpha-1)+m-1}{2\sqrt{x}}\mathbf{B}_{m}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}$
@@
where $\mathbf{B}_{m}$ is given as in (A.3), $\mathbf{C}^{(1)}_{m}=\big(c_{j,k}^{(1)}\big)_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-1\end{subarray}}$ is an $m\times(m-1)$ matrix satisfying
$c_{j,k}^{(1)}=\begin{cases}(-1)^{j+k}\bigl(\frac{m-1-k}{k+1}+\frac{2l-k+\alpha}{k(k+1)}\bigr),&\text{if }j\leq k\leq m-1,\\ -\frac{1}{k+1}(k(2l-k+\alpha)+m-1-k),&\text{if }k=j-1,\\ 0,&\text{if }k<j-1,\end{cases}$
and $\mathbf{C}^{(2)}_{m}=\bigl(c_{j,k}^{(2)}\bigr)_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-2\end{subarray}}$ is an $m\times(m-2)$ matrix satisfying
$c_{j,k}^{(2)}=\begin{cases}(-1)^{j+k}\frac{l+2}{(k+1)(k+2)},&j-1\leq k\leq m-2,\\ \frac{j-l-2}{j},&k=j-2,\\ 0,&k<j-2.\end{cases}$
@@
The above proposition generalizes [39, Proposition 4.3] from the case $\alpha=1$ to general $\alpha$. It can be rewritten as a first-order matrix linear differential equation for the vector on the left-hand side analogous to (but distinct from) that implied by Proposition 3.3, with $n=m$, $\beta=2$ and $B_{p}$ given by (3.13) with $q=m+\alpha$. On this latter point, note that the entries in the corresponding vectors agree in the final entry but involve different functions otherwise.
Theorem A.8.
@@
$\tau_{l,Y_{l+1,l+1},\alpha}(x)=x^{-\frac{l+1}{2}}\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{[\frac{l+1-n}{2}]}b_{j,n}^{(l+1)}(\alpha)x^{j}\right)\frac{{\rm d}^{n}\tau_{l,\alpha}(x)}{{\rm d}x^{n}},$ (A.9)
where the coefficients $b_{j,n}^{(l+1)}(\alpha)$ are given recursively. Moreover, we have the following equation:
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{[\frac{l+1-n}{2}]}b_{j,n}^{(l+1)}(\alpha)x^{j}\right)\frac{{\rm d}^{n}\tau_{l,\alpha}(x)}{{\rm d}x^{n}}=0.$ (A.10)
@@
$\tau_{l,Y_{l+1,l+1},\alpha}(x)={}$ $\frac{1}{l+1}\left(\sqrt{x}\frac{{\rm d}\tau_{l,Y_{l,l},\alpha}(x)}{{\rm d}x}-\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)\right)+\frac{l(l+\alpha)}{(l+1)\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)$
@@
$\tau_{l,Y_{l-1,l-1},\alpha}(x)=\operatorname{Tr}\bigl(\mathrm{adj}\bigl(\big[I_{j+k+\alpha+1}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}\bigr)\big[I_{j+k+\alpha}\bigl(2\sqrt{x}\bigr)\big]_{j,k=0}^{l-1}\bigr).$
By the observation $\tau_{l,Y_{l,l},\alpha}(x)=\det\bigl(I_{j+k+\alpha+1}\bigl(2\sqrt{x}\bigr)\bigr)_{j,k=0,\ldots,l-1}$ and the recurrence relation
$I_{j+k+\alpha}\bigl(2\sqrt{x}\bigr)=\sqrt{x}\frac{\rm d}{{\rm d}x}I_{j+k+\alpha+1}\bigl(2\sqrt{x}\bigr)+\frac{\alpha+j+k+1}{2\sqrt{x}}I_{j+k+\alpha+1}\bigl(2\sqrt{x}\bigr),$
@@
$\sqrt{x}\frac{\rm d}{{\rm d}x}\tau_{l,Y_{l,l},\alpha}(x)+\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)-\tau_{l,Y_{l-1,l-1},\alpha}(x)=0.$
Then by (A.11), we have the identity $\tau_{l,Y_{l+1,l+1},\alpha}\equiv 0$. Hence, using above and (A.9), we obtain (A.10). ∎
@@
As a consequence of Theorem A.8 with $\alpha=1$, we obtain the following result, which provides a method to derive the degree $l+1$ linear differential equation satisfied by (1.7).
@@
Let $l\geq 1$. Let $b_{j,n}^{(l+1)}(1)$ be given as in Theorem A.8 with $\alpha=1$. Then
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{[\frac{l+1-n}{2}]}b_{j,n}^{(l+1)}(1)x^{j}\right)\frac{{\rm d}^{n}\tau_{l}(x)}{{\rm d}x^{n}}=0.$
Finally, applying Corollary A.10 and verifying with the code, [^5] https://drive.google.com/file/d/1gd4kQO9ABKhALrzeYos-DbeLcelQmik6/view?usp=share_link we re-obtain Proposition 4.6.
Appendix B Computational aspects. Appendix by Folkmar Bornemann∗*∗**Department of Mathematics, Technical University of Munich, 80290 München, GermanyE-mail: bornemann@tum.de
B.1 $\boldsymbol{D}$-finiteness